OpenStudy (anonymous):

find the solution to the equation 5|2x-5|=5

OpenStudy (anonymous):

|2x-5|=1 2x-5=1 and 2x-5=-1 .'. x=3 and x=2

OpenStudy (anonymous):

(3,2)

OpenStudy (anonymous):

Here is the equation in wolfram Alpha - just in case you wanted to see it plotted. http://www.wolframalpha.com/input/?i=solve+5 |2x-5|%3D5+

OpenStudy (anonymous):

thanks